OpenStudy (anonymous):

Im having massive brain block. -exp(t) + 3exp(t/2) = 0, solve for t

OpenStudy (anonymous):

replace \(\frac{t}{2}\) by \(x\) and solve \[-e^{2x}+3e^x=0\]

OpenStudy (anonymous):

you move one term to the other side and take natural log of both sides?

OpenStudy (anonymous):

solve as a quadratic equation

OpenStudy (anonymous):

factor out an \(e^x\)

OpenStudy (anonymous):

\[-e^x(e^x-3)=0\]

OpenStudy (anonymous):

exp(x) = 3 or 0 but never gets to 0

OpenStudy (anonymous):

right

OpenStudy (anonymous):

make sure to replace \(x\) by \(\frac{t}{2}\)

OpenStudy (anonymous):

ok, thank you very much!

OpenStudy (anonymous):

yw